OpenStudy (anonymous):

Can 32/160 be simplified?

OpenStudy (anonymous):

\[\frac{ 32 }{160 }\] Yes it can be simplified/reduced. You can divide 8 into both of them making it \[\frac{ 4 }{20 }\] which you can reduce this new fraction dividing both by 4 to make the answer: \[\frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

Thanks (: